OpenStudy (awkwardpanda):

Rewrite each equation for each line in slope intercept form. Graph each line. 1. -4+6y=18 2. 3y+3x=15 3. 3y+x=5

OpenStudy (wolfe8):

Write into the general form y=mx+b Can you do that?

OpenStudy (shamil98):

panda we went over this.

OpenStudy (awkwardpanda):

IM SOWWY.

OpenStudy (shamil98):

-.-

OpenStudy (shamil98):

to write those equations in slope-intercept form. YOU MUST GET Y BY ITSELF. -4x + 6y = 18 add 4x to both sides 6y = 4x + 18 and divide both sides by 6. 6y/6=4x/6 + 18/6 y= (2/3)x +3

OpenStudy (shamil98):

3 is the y intercept you start from , then you graph the line by the slope (2/3) so up two and right 3, and you do the opposite to get the bottom part of the line so you would make the slope -(2/3) so down two and left 3.

OpenStudy (shamil98):

do ya understand

OpenStudy (awkwardpanda):

yES ;-;

OpenStudy (shamil98):

To make sure, putting the second one in slope-intercept form.

OpenStudy (shamil98):

try*

OpenStudy (awkwardpanda):

Okay ;-;

OpenStudy (shamil98):

remember slope-intercept form is y = mx +b

OpenStudy (awkwardpanda):

Thank you Sham ;-;

OpenStudy (awkwardpanda):

WHATS 2. HALP. IDK IM SOWWY. @shamil98

OpenStudy (shamil98):

I KNEW IT

OpenStudy (shamil98):

TRY DOING IT

OpenStudy (awkwardpanda):

;-;

OpenStudy (shamil98):

u can do uht

OpenStudy (awkwardpanda):

so i divide by 3y? ;-;

OpenStudy (shamil98):

first you need to get y by itself there is a positive 3x, so to get it to the other side...

OpenStudy (shamil98):

you want to get rid of it in math it is called ..... subtraction

OpenStudy (awkwardpanda):

So.. subtract 3y ;-;

OpenStudy (shamil98):

nooo.. substract 3x. ..

OpenStudy (awkwardpanda):

OH. I WAS LOOKING AT THE WRONG ONE ;-;

OpenStudy (shamil98):

in the 3d one you still would subtract x ._.

OpenStudy (awkwardpanda):

aHem.. 3x=3y+15 divide by 3..?

OpenStudy (shamil98):

noo, you are solving for Y

OpenStudy (shamil98):

y = mx + b , thats slope-intercept ._.

OpenStudy (awkwardpanda):

so y=3x+15..

OpenStudy (shamil98):

3y = -3x + 15

OpenStudy (shamil98):

you subtract 3x from both sides .

OpenStudy (awkwardpanda):

gosh flab it. okay so it would be

OpenStudy (shamil98):

now you divide by 3

OpenStudy (shamil98):

you want to get y just by itself .

OpenStudy (shamil98):

3y /3 = y -3x/3 = 15/3 =

OpenStudy (awkwardpanda):

So step one is 3y+3x=15 step 2 is dividing 3x which would make it 3y= -3x+15 step 3 is dividing by 3 step 4 is y=x+5

OpenStudy (shamil98):

step 2 is subtracting 3x * step 4 is y = -x +5* remember its a NEGATIVE 3x so it would -x :)

OpenStudy (shamil98):

but you got most of it , so its a start.

OpenStudy (awkwardpanda):

yay! okay im gonna start again

OpenStudy (shamil98):

try number 3.

OpenStudy (shamil98):

3y + x = 5 what's the first step? :3

OpenStudy (awkwardpanda):

 just to make it clear XD and for that one it's divide by x?